OpenStudy (anonymous):

Calculus II - Volume By Shells Hello, having some problems with a volume by shells problem. I believe my issue is the setup. Question: Use the shells method to find the volume of a solid generated by revolving the region bounded by the lines y=x+6 and y=x^2 about the following lines; a) The line x=3 b) The line x=-2 c) The x-axis d) The line y=9 So - I convert; y = x+6 x = y-6 y = x^2 x = sqrt(y) Rest coming in an equation post -

OpenStudy (anonymous):

\[2\pi \int\limits_{0}^{9} (3-x)(\sqrt(y) - (y-6))\] Is this the correct setup?

OpenStudy (turingtest):

for a ?

OpenStudy (anonymous):

Yeah, sorry.

OpenStudy (turingtest):

well I would set it up the other way, in terms of x...

OpenStudy (anonymous):

That was a typo, I meant to have (3-y)(sqrt(y) etc, etc...

OpenStudy (turingtest):

\[2\pi\int_{-2}^{3}(3-x)[(x+6)-x^2]xdx\]is the way I see it easiest...